The Wagner Company currently uses an injection-molding machine that was purchased 2 years ago. This machine is being depreciated on a straight-line basis, and it has 6 years of remaining life. Its current book value is $2,100, and it can be sold for $2,500 at this time. Thus, the annual depreciation expense is $2,100/6=$350 per year. If the old machine is not replaced, it can be sold for $500 at the end of its useful life. Wagner is offered a replacement machine that has a cost of $8,000, an estimated useful life of 6 years, and an estimated salvage value of $800. This machine falls into the MACRS 5-year class, so the applicable depreciation rates are 20%, 32%, 19%, 12%, 11%, and 6%. The replacement machine would permit an output expansion, so sales would rise by $1,000 per year; even so, the new machine’s much greater efficiency would reduce operating expenses by $1,500 per year. The new machine would require that inventories be increased by $2,000, but accounts payable would simultaneously increase by $500. Wagner’s marginal federal-plus-state tax rate is 40%, and its WACC is 15%. Should it replace the old machine?

Please do not do the calculations in excel, I am trying to follow all of the steps through for setting up the NPV equation.
Answer #1
DEPRECIATION OF OLD MACHINE
Year from today 1 2 3 4 5 6
D1 Annual Depreciation expense $350 $350 $350 $350 $350 $350
Salvage Value $500
DEPRECIATION OF NEW MACHINE
Year(from today) 1 2 3 4 5 6
A Depreciation Rate 20% 32% 19% 12% 11% 6%
D2=A*$8000 Depreciation amount $1,600 $2,560 $1,520 $960 $880 $480
Present value of Cash Flow=(Cash Flow)/((1+i)^N)
i=discount Rate =WACC =15%=0.15
N=Year of Cash Flow
N Year From Today 1 2 3 4 5 6
D=D2-D1 Incremental Depreciation $1,250 $2,210 $1,170 $610 $530 $130
E=D*40% Depreciation tax shield $500 $884 $468 $244 $212 $52
B Before tax saving in operating cost $1,500 $1,500 $1,500 $1,500 $1,500 $1,500
C Increase in Sales $1,000 $1,000 $1,000 $1,000 $1,000 $1,000
F=B+C Before tax increase in operating Cash flow $2,500 $2,500 $2,500 $2,500 $2,500 $2,500
G=F*(1-0.4) After tax increase in operating cash flow $1,500 $1,500 $1,500 $1,500 $1,500 $1,500
TerminalCash Flows;
Before tax incremental Salvage Value(800-500) $300
H After tax Incremental Salvage value =300*(1-0.4) $180
J Release of additional working capital(2000-500) $1,500
K=H+J Total TerminalCash Flow $1,680
L=E+G+K Total Cash Inflow $2,000 $2,384 $1,968 $1,744 $1,712 $3,232 SUM
M=L/(1.15^N) Present Value of Cash Inflow $1,739 $1,803 $1,294 $997 $851 $1,397 $8,081
PV Sumof Present Value of Cash inflows $8,081
Initial Outlay:
Cost of new machine $8,000
Salvage Value of old machine ($2,500)
Increase in working capital $1,500 (2000-500)
I Initial Outlay: $7,000
NPV=PV-I Net Present value $1,081 (8081-7000)
The Machine Should be replace
NPV is Positive